OpenStudy (1dfanatic):

Find the range of the function s(t) = t^2 - 1 given the domain { -1, 0.5, 3.7 }

OpenStudy (solomonzelman):

You know that, \(\color{#000000 }{ \displaystyle t^2 }\) is at least \(\color{#000000 }{ \displaystyle 0 }\). (i.e. \(\color{#000000 }{ \displaystyle t^2\ge0 }\))

OpenStudy (solomonzelman):

You can subtract 1 from both sides and this will give you the domain.

OpenStudy (1dfanatic):

Im sorry i still don't get it

OpenStudy (solomonzelman):

Oh, my bad, you have the given ranges, I didn't see them.

OpenStudy (solomonzelman):

For any value in the domain, you need to plug in this value, to find the output.

OpenStudy (solomonzelman):

The inputs = domain, The (corresponding) outputs = range.

OpenStudy (solomonzelman):

So, start from plugging \(\color{#000000 }{ t=-1 }\), into \(\color{#000000 }{ \displaystyle s(t)=t^2-1 }\).

OpenStudy (1dfanatic):

I got (t+1)(t-1)...

OpenStudy (solomonzelman):

If \(\color{#000000 }{ \displaystyle t=-1 }\), then \(\color{#000000 }{ \displaystyle s(-1)=(-1)^2-1 =1-1=0}\).

OpenStudy (1dfanatic):

I'm still so confused, sorry..